Michael and Jackson are standing on opposite corners of a ball field that is 60 yards long and 30 yards wide. Michael walks on the edge of the field for one length and one width. Jackson walks across the diagonal of the field. How many more yards did Michael walk than Jackson?
a
around 157 yards more
b
around 23 yards more
c
90 yards more
d
Michael didn’t walk farther, Jackson did

Answer by Boreal(15235)

Pythagorean theorem for diagonal.
60^2+30^2=d^2=4500
d=sqrt(4500)=67.1 yds. Jackson's distance
Michael walked 60+30 or 90 yds., or 23 yen more.
b.